The area of the wall was 78 square feet. What was the length of one side of the wall?
Answer by Alan3354(69443)

The area of the wall was 78 square feet. What was the length of one side of the wall?
---------------
Area = L*W
78 = L*W
L = 78/W
------------
That's the length of both sides, if it's a rectangle (most walls are).